I've got 3 that I'm stuck on out of 25...............
Solve for x
4x + 16 =20

1/2x - 4 =12

3x + 5 = x + 17
Frustration has set in or i'm just too tired!

Answer by elima(1433)

1) 4x + 16 =20
First you will subtract 126 from both sides:
4x+16-16=20-16
4x=4
Now divide 4 from both sides:
4x/4=4/4
x=1
2)1/2x - 4 =12
First add 4 to both sides:
1/2x-4+4=12+4
1/2x=16
Then you will multiply the inverse of 1/2(which is2/1) to both sides:
1/2x(2/1)=16(2/1)
x=32
3) 3x + 5 = x + 17
First subtract 5 from both sides;
3x+5-5=x+17-5
3x=x+12
Now subtract x from both sides;
3x-x=x+12-x
2x=12
Now divide both sides by 2;
2x/2=12/2
x=6
Hope this helps, I now when your tired nothing makes sense =)